Opened 9 years ago

Closed 8 years ago

Last modified 3 years ago

#361 closed flight processing (fixed)

EUFAR10/07 (ADDRESSS), flight day 234b, Summer School Geo Site

Reported by: iopa Owned by: knpa
Priority: alpha 4 high Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors: emca,anch

Description (last modified by knpa)

Data location: ~arsf/workspace/EUFAR10_07-2010_234b_Summer_School_Geo_Site

Data arrived from ARSF via SATA disk Q on 15/09/2010.

Scientific objective: Mapping of lithology, fault structure,
erosion parameters and soil components

Priority: a4h

PI: A. Zlinszky

Note: Data is currently symlinked from ~arsf/arsf_data/2010/flight_data/archive_overflow and when this is resolved the symlinks will need to be re-directed.

PI has requested that data be made available via NEODC promptly as well as normal delivery - data will be downloaded from NEODC by students.

Sensors:

  • Eagle (04/03/2011)
  • Hawk (04/03/2011)
  • Leica LIDAR (20/04/2011)
  • Digital Photography (04/03/2011)

Change History (35)

comment:1 Changed 9 years ago by iopa

  • Summary changed from EUFAR10/07 to EUFAR10/07, flight day 234b, Summer School Geo Site

comment:2 Changed 9 years ago by iopa

  • Description modified (diff)

comment:3 Changed 9 years ago by iopa

Applanix and IPAS have been done. The applanix data currently in the workspace covers both this project and also 234c.
For IPAS, there is processed IPAS in 234a in the workspace that covers 234a,234b,234c.
The proc directory from there could be entirely copied over to 234b or just symlinked, your choice.

comment:4 Changed 9 years ago by knpa

  • Description modified (diff)

Added scientific objective.

comment:5 Changed 9 years ago by benj

  • Description modified (diff)

comment:6 Changed 9 years ago by benj

  • Description modified (diff)

PI has requested that data be made available via NEODC promptly as well as normal delivery - data will be downloaded from NEODC by students.

comment:7 Changed 9 years ago by anch

Starting hyperspectral processing

comment:8 Changed 9 years ago by emca

Processed the camera imagery from raw to tiff format

comment:9 Changed 9 years ago by anch

Hyperspectral ready for delivery check at:

/users/rsg/arsf/workspace/EUFAR10_07-2010_234b_Summer_School_Geo_Site/delivery/20110218/EUFAR10_07

comment:10 Changed 9 years ago by anch

Eagle and Hawk line 5 (006 on logsheet) was not recorded correctly. Images were still captured over the flightline but the hyperspectral data captured was all over a turn where the plane was getting ready for the next flightline and can't be processed using azgcorr

comment:11 Changed 9 years ago by knpa

  • Other processors set to emca,anch
  • Owner set to knpa
  • Status changed from new to assigned

Taking LIDAR.

comment:12 Changed 9 years ago by iopa

Checking hyperspectral.

comment:13 Changed 9 years ago by iopa

Hyperspectral delivery check;

  • need to include flight letter in lev1's and screenshots

all else seems good and ready to deliver.

comment:14 Changed 9 years ago by anch

Filenames fixed, and readme updated to reflect this.

comment:15 Changed 9 years ago by knpa

Processed LIDAR lines.

All have acceptable vertical and horizontal discrepancies except line 1 (09:26) which had a large roll error. Corrected with roll value 0.00125. All others keep default values.

comment:16 Changed 9 years ago by emca

Starting camera processing

comment:17 Changed 9 years ago by anch

Finished camera delivery for emca as she's on leave.

Have kept a copy of Readme.tex as some manual editing was required due complications with line 006 (As detailed above).

Delivery is now ready for checking here:

/users/rsg/arsf/workspace/EUFAR10_07-2010_234b_Summer_School_Geo_Site/delivery/20110303/EUFAR10_07

comment:18 Changed 9 years ago by iopa

Checking camera.

comment:19 Changed 9 years ago by iopa

Camera check;

  • as an observation, flightline 5 in the KML is partly missing since there is no eagle/hawk data as commented previously in this ticket.

otherwise data and delivery are good to go.

comment:20 Changed 9 years ago by anch

  • Description modified (diff)

Camera sent to PI

comment:21 Changed 9 years ago by anch

  • Description modified (diff)

Hyperspectral was also sent to PI 04/03/2011

comment:22 Changed 9 years ago by anch

  • Summary changed from EUFAR10/07, flight day 234b, Summer School Geo Site to EUFAR10/07 (ADDRESSS), flight day 234b, Summer School Geo Site

comment:23 Changed 9 years ago by emca

Delivery checked the LiDAR delivery at
~airborne/workspace/EUFAR10_07-2010_234b_Summer_School_Geo_Site/delivery/20110408

All data looks good. Just need to include the LAS files with the delivery.

comment:24 Changed 9 years ago by knpa

LAS files included with lidar delivery as per above.

LIDAR copied on to hard disk 48. 238c LIDAR is awaiting delivery check so will wait to put that on disk also.

comment:25 Changed 9 years ago by emca

  • Component changed from Processing: general to Archiving

All data has been rsync'd to
/users/rsg/arsf/arsf_data/2010/flight_data/archive-overflow/EUFAR10_07-2010_234b_Summer_School_Geo_Site

Ready for archiving

comment:26 Changed 9 years ago by emca

LAS files included with delivery so removing from leica/proc_laser for archiving

Aimms folder is present, but contained file is empty - not going to archive

Grimm and tabi data present

comment:27 Changed 9 years ago by emca

Leica IPAS project files are missing, have the processed files

comment:28 Changed 9 years ago by emca

Archive tarball created and made available to NEODC

comment:29 Changed 9 years ago by knpa

  • Description modified (diff)

LIDAR delivered today to A. Zlinszky in Tihany, Hungary (20/04/11)

comment:30 Changed 9 years ago by knpa

Backed up by NEODC, moved to archive disk.

comment:31 Changed 8 years ago by knpa

  • Milestone set to 2010 data processing completion
  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace version. Closing ticket.

comment:32 Changed 7 years ago by adbe

Hyperspectral Delivery Check:

  • Mention dropped scans in ReadMe

comment:33 Changed 7 years ago by emca

This data was processed by jaho - updated the Readme for dropped scans.

Reprocessing now complete. Adding this project to rest of the data being sent to A Zlinszky.

Last edited 7 years ago by emca (previous) (diff)

comment:34 Changed 7 years ago by emca

Data sent to A Zlinzky 16/08/2012

Note: See TracTickets for help on using tickets.